Calculate pH for each H3O+ concentration:
Answer #1

1) pH = -log[H3O+] = -log(1*10-7 M) = 7

2) pH = -log[H3O+] = -log(1*10-3 M) = 3

3)  pH = -log[H3O+] = -log(1*10-12 M) = 12